OpenStudy (anonymous):

All real numbers between -4 and 1, inclusive

OpenStudy (anonymous):

\[-4\le x \le 1 or \left[ -4,1 \right]\]

OpenStudy (zzr0ck3r):

There are many of them. I fact, there are just as many numbers between -4 and 1, as there are numbers period. In fact, there are just as many numbers between 0.0000001 and 0.00000001 as there are numbers!